OpenStudy (anonymous):

i'm new on this website and i need help on finding: cos inverse[cos(4 pi/5)] and earning medals and such! (^-^)

OpenStudy (anonymous):

THE ANSWER IS 2.51

OpenStudy (anonymous):

\[\cos^{-1}(\cos (4 \pi / 5)) = 4 \pi / 5\] = 2.513274123 cos inverse cos cancel each other